OpenStudy (anonymous):

cos 70◦ + sin 80◦ = ?

OpenStudy (unklerhaukus):

change that cosine into a sine, \[\cos(90°-u)=\sin (u)\]

OpenStudy (unklerhaukus):

then use the \[\sin u+\sin v=2\sin\Big(\frac{u+v}2\Big)\cos\Big(\frac{u-v}2\Big)\]

OpenStudy (anonymous):

the choices are a.\[\sqrt{3} \sin 50◦\] b.\[\sqrt{3} \cos 50◦\] c.\[\sqrt{3} \sin 40◦\] d.\[\sqrt{3} \cos 10◦\] e.\[\sqrt{3} \sin 10◦\]

OpenStudy (anonymous):

can you help me, what is the correct answer?

OpenStudy (anonymous):

what is the correct answer?

OpenStudy (unklerhaukus):

or maybe it would have been easier to change the sine to a cosine \[\sin(90°-u)=\cos (u)\] \[\sin(80°)=\sin (90°-10°)=\cos(10°)\]

OpenStudy (unklerhaukus):

\[\cos 70° + \sin 80°\\ =\cos 70° + \cos10° \] then use \[\cos(u)+\cos(v)=2\cos\big(\frac{u+v}2\big)\cos\big(\frac{u-v}2\big)\]